Jake is five years older than laura. Four years ago jake was twice as old as laura was then. How old is jake now?
Answer by lwsshak3(11628)

Jake is five years older than laura. Four years ago jake was twice as old as laura was then. How old is jake now?
**
let x= jake's age now
x-5=laura's age now
four years ago:
jake's age=x-4
laura's age=x-5-4=x-9
(x-4)=2(x-9)
x-4=2x-18
-x=-14
x= 14
jake's age now=14